Whats the complete solution of this linear equation?
2x + y − z = −12
−x + y + z = 7
−2x + 4z = 34
=

Answer by stanbon(75887)

Whats the complete solution of this linear equation?
2x + y − z = −12
−x + y + z = 7
−2x + 4z = 34
--------------
I used a matrix method on my TI-84 to get:
x = -1
y = -2
z = 8
==============
Cheers,
Stan H.
-----------